### v2

# Non-uniqueness of positive solutions for supercritical semilinear heat equations without scale invariance

Abstract.
We establish nonuniqueness of solutions for Cauchy problems of semilinear heat equations with a wide class of nonlinearities. Specifically, we consider
$\begin{cases}\partial_{t}u-\Delta u=f(u),&x\in\mathbb{R}^{N},\ t>0,\\ u(x,0)=u_{0}(x),&x\in\mathbb{R}^{N},\end{cases}$
where $N>2$. We assume that the growth rate of $f$ is less than the Joseph-Lundgren exponent for $N>10$ and it satisfies certain assumptions guaranteeing the existence of a positive radial singular stationary solution $u^{*}$. We prove that if $u_{0}=u^{*}$, then the problem has at least two positive solutions, namely $u^{*}$ and $u(t)$ which satisfies $u(t)\in L_{loc}^{\infty}(0,t_{0};L^{\infty}(\mathbb{R}^{N}))$ for some $t_{0}>0$ and
$u(t)\to u^{*}\quad\text{in}\ L^{\gamma}_{ul}(\mathbb{R}^{N})\quad\text{as}\ t\to 0^{+}$
for $1\leq\gamma<N(p_{f}-1)/2$, where $p_{f}:=\lim_{u\to\infty}uf^{\prime}(u)/f(u)$ is a growth rate of $f$. Hence, nonuniqueness problem can be reduced to the existence problem of a positive radial singular stationary solution. The method of construction of $u(t)$ is based on the monotonicity argument. Transformations of forward self-similar solutions for $f(u)=u^{p}$ and $e^{u}$ play a crucial role.
Key words and phrases:
Self-similar solutions; Singular stationary solution; Joseph-Lundgren exponent; Quasi-scaling
2020 Mathematics Subject Classification:
Primary: 35K15, 35A02, secondary 35A21, 35B44.
ORCiD of YM is 0000-0002-7766-1849
YM was supported by JSPS KAKENHI Grant Number 24K00530.
1. Introduction and main results
We are concerned with Cauchy problem for the semilinear heat equation
(1.1) $\left\{\begin{aligned} &\partial_{t}u-\Delta u=f(u),\quad&&x\in\mathbb{R}^{N},\,\,t>0,\\ &u(x,0)=u_{0}(x),\quad&&x\in\mathbb{R}^{N},\end{aligned}\right$.
where $N>2$ and $u_{0}$ is a nonnegative measurable function on $\mathbb{R}^{N}$. The aim of this paper is to establish nonuniqueness of positive solutions of problem (1.1) for a wide class of nonlinearities if problem (1.1) has a positive radial singular stationary solution.
Let us explain notations. Let $p_{S}$ denote the critical Sobolev exponent, i.e.,
$p_{S}:=\begin{cases}\displaystyle{\frac{N+2}{N-2}}&\text{if}\ N>2,\\ \infty&\text{if}\ N=1,2.\end{cases}$
Throughout the present paper, we make the following assumptions on $f$:
  • (A1)
    $f\in C^{1}[0,\infty)\cap C^{2}(0,\infty)$, $f(0)=0$, and $f^{\prime}(0)=0$;
  • (A2)
    $f^{\prime}(u)>0$ for $u>0$ and $f^{\prime\prime}(u)>0$ for $u>0$;
  • (A3)
    There exists a limit
    $q_{f}:=\lim_{u\to\infty}\frac{f^{\prime}(u)^{2}}{f(u)f^{\prime\prime}(u)};$
  • (A4)
    The following holds:
    $Q(u):=uf(u)-(p_{S}+1)\int_{0}^{u}f(s)\,\mathop{}\!\mathrm{d}s\geq 0\quad\mbox{for}\ u\geq 0.$
We define $f(u)=0$ for $u<0$. In [8] $A$ is defined by $A=\lim_{u\to\infty}f^{\prime}(u)F(u)$, where
(1.2) $F(u)=\int_{u}^{\infty}\frac{ds}{f(s)}\quad\mbox{for}\quad u>0.$
Then, when (A3) holds, by L’Hospital rule we have
$A=\lim_{u\to\infty}\frac{F(u)}{1/f^{\prime}(u)}=\lim_{u\to\infty}\frac{f^{\prime}(u)^{2}}{f(u)f^{\prime\prime}(u)}=q_{f}.$
It is proved in [8, Remark 1.1] that if the limit $A$ exists, then $A\geq 1$. Hence, $q_{f}\geq 1$ also holds if (A3) holds. The growth rate of $f$ can be defined by
$p_{f}:=\lim_{u\to\infty}\frac{uf^{\prime}(u)}{f(u)}.$
Then, by L’Hospital rule we have
$\frac{1}{p_{f}}=\lim_{u\to\infty}\frac{f(u)/f^{\prime}(u)}{u}=\lim_{u\to\infty}\left(1-\frac{f(u)f^{\prime\prime}(u)}{f^{\prime}(u)^{2}}\right)=1-\frac{1}{q_{f}},$
and hence $1/p_{f}+1/q_{f}=1$. The exponent $q_{f}$ is the conjugate exponent of the growth rate $p_{f}$. Using the $q_{f}$-exponent, we can treat power nonlinearities ($1<q_{f}<\infty$) and super-power nonlinearities ($q_{f}=1$), including exponential nonlinearities, in a unified way. We later use the conjugate exponent of $p_{S}$, i.e.,
$q_{S}:=\frac{p_{S}}{p_{S}-1}=\frac{N+2}{4}.$
The exponent $q_{S}$ can be formally defined for $N\geq 1$. However, $q_{S}$ becomes effective only in the case $N>2$, since $q_{S}>1$ for $N>2$ and $q_{f}\geq 1$ always holds.
Typical examples are $f(u)=u^{\beta}$, $\beta\geq p_{S}$, and $f(u)=u^{\beta}e^{u^{\gamma}}$, $\beta\geq p_{S}$ and $\gamma>1$. Other examples are given in Section 2.
We need the so-called Joseph-Lundgren exponent
$p_{JL}:=\begin{cases}\displaystyle{1+\frac{4}{N-4-2\sqrt{N-1}}}&\text{if}\ N>10,\\ \infty&\text{if}\ 2\leq N\leq 10.\end{cases}$
The conjugate exponent of $p_{JL}$ is given by
$q_{JL}:=\frac{p_{JL}}{p_{JL}-1}=\frac{N-2\sqrt{N-1}}{4}.$
The exponent $q_{JL}$ can be formally defined for $N\geq 1$. However, $q_{JL}$ becomes effective only in the case $N>10$, since $q_{JL}>1$ for $N>10$ and $q_{f}\geq 1$ always holds.
For $x\in\mathbb{R}^{N}$ and $\rho>0$, set $B(x,\rho):=\left\{y\in\mathbb{R}^{N};\ |x-y|<\rho\right\}$. We define a uniformly local $L^{p}$ space as follows: For $1\leq p\leq\infty$,
$L^{p}_{{ul}}(\mathbb{R}^{N}):=\left\{u\in L^{p}_{{loc}}(\mathbb{R}^{N});\ \left\|u\right\|_{L^{p}_{ul}(\mathbb{R}^{N})}<\infty\right\},$
where
$\left\|u\right\|_{L^{p}_{{ul}}(\mathbb{R}^{N})}:=\begin{cases}\displaystyle{\operatorname{ess\,sup}\displaylimits_{z\in\mathbb{R}^{N}}\left\|u\right\|_{L^{\infty}(B(z,1))}}&\textrm{if}\ p=\infty,\\ \displaystyle{\sup_{z\in\mathbb{R}^{N}}\left(\int_{B(z,1)}|u(x)|^{p}\,\mathop{}\!\mathrm{d}x\right)^{\frac{1}{p}}}&\textrm{if}\ 1\leq p<\infty.\end{cases}$
It is obvious that $L^{\infty}_{{ul}}(\mathbb{R}^{N})=L^{\infty}(\mathbb{R}^{N})$ and that $L^{p_{1}}_{{ul}}(\mathbb{R}^{N})\subset L^{p_{2}}_{{ul}}(\mathbb{R}^{N})$ if $1\leq p_{2}\leq p_{1}<\infty$. Let $\mathcal{L}^{p}_{{ul}}(\mathbb{R}^{N})$ denote the closure of the space of bounded uniformly continuous functions $BUC(\mathbb{R}^{N})$ in the space $L^{p}_{{ul}}(\mathbb{R}^{N})$, i.e.,
(1.3) $\mathcal{L}^{p}_{{ul}}(\mathbb{R}^{N}):=\overline{BUC(\mathbb{R}^{N})}^{\left\|\,\cdot\,\right\|_{L^{p}_{{ul}}(\mathbb{R}^{N})}}.$
Let $G_{t}=G_{t}(x)$ be the fundamental solution of
$\partial_{t}v-\Delta v=0\quad\mbox{in}\quad\mathbb{R}^{N}\times(0,\infty),$
that is,
$G_{t}(x):=\frac{1}{(4\pi t)^{\frac{N}{2}}}\exp\left(-\frac{|x|^{2}}{4t}\right)$
for $x\in\mathbb{R}^{N}$ and $t>0$. For $u_{0}\in L^{1}_{{ul}}(\mathbb{R}^{N})$, let $S(t)u_{0}$ be
$S(t)u_{0}](x):=\int_{\mathbb{R}^{N}}G_{t}(x-y)u_{0}(y)\,\mathop{}\!\mathrm{d}y$
for $x\in\mathbb{R}^{N}$ and $t>0$.
We are in a position to define a solution of problem (1.1).
Definition 1.1.
Let $u$ be a nonnegative measurable function on $\mathbb{R}^{N}\times[0,T)$, where $T\in(0,\infty]$.
  • (i)
    We say that $u$ is a solution of problem (1.1) in $\mathbb{R}^{N}\times[0,T)$ if
    (1.4) $\begin{split}\infty>u(x,t)&=\int_{\mathbb{R}^{N}}G_{t}(x-y)u_{0}(y)\,\mathop{}\!\mathrm{d}y+\int_{0}^{t}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)f(u(y,s))\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\end{split}$
    for a.a. $(x,t)\in\mathbb{R}^{N}\times[0,T)$.
  • (ii)
    If $u$ satisfies (1.4) with == replaced by $\geq$, then we say that $u$ is a supersolution in $\mathbb{R}^{N}\times[0,T)$.
Under the assumptions (A1)–(A4) with $q_{f}<q_{S}$, the following positive radial singular stationary solution exists:
Proposition 1.2.
Let $N>2$. Suppose (A1)(A4). Suppose, in addition, that $q_{f}<q_{S}$. The problem
(1.5) $\begin{cases}-\Delta u=f(u),&x\in\mathbb{R}^{N}\setminus\{0\},\\ u(x)\geq 0,&x\in\mathbb{R}^{N}\setminus\{0\},\\ u(x)\ \mbox{is radial},\\ u(x)\to\infty\quad\mbox{as}\quad|x|\to 0^{+},\end{cases}$
has a unique solution $u^{*}\in C^{2}(\mathbb{R}^{N}\setminus\{0\})$, which is called a singular solution. Moreover, $u^{*}\in\mathcal{L}^{\gamma}_{{ul}}(\mathbb{R}^{N})$ for $1\leq\gamma<\gamma^{*}$, $u^{*\prime}(r)<0$ for all $r>0$ and $u^{*}$ satisfies
$u^{*}(x)=F^{-1}\left[\frac{|x|^{2}}{2N-4q_{f}}(1+o(1))\right]\quad\mbox{as}\quad|x|\to 0^{+},$
where
[…]
and $F^{-1}$ is the inverse function of $F$ defined by (1.2).
We will prove the following lemma in Section 3.
Lemma 1.3.
Let $u^{*}$ be as in Proposition 1.2. Then, $u^{*}$ is a positive singular stationary solution of problem (1.1) in the sense of Definition 1.1. In particular,
$\infty>u^{*}(x)=[S(t)u^{*}](x)+\int_{0}^{t}[S(t-s)f(u^{*})](x)\,\mathop{}\!\mathrm{d}s$
for all $x\in\mathbb{R}^{N}\setminus\{0\}$ and $t\in(0,\infty)$.
The main result of the present paper is the following:
Theorem 1.4 (Nonuniqueness (supercritical)).
Let $N>2$. Suppose (A1)(A4). Let $u^{*}$ be the singular solution of problem (1.5) given in Proposition 1.2. Suppose, in addition, the following (A5) and (A6):
  • (A5)
    $q_{JL}<q_{f}<q_{S}\quad$ (“almost” equivalently, $p_{S}<p_{f}<p_{JL}$);
  • (A6)
    If $q_{f}=1$, then $\frac{f^{\prime}(u)^{2}}{f(u)f^{\prime\prime}(u)}\leq 1$ for large $u>0$.
Then, problem (1.1) with $u_{0}=u^{*}$ has at least two positive solutions in the sense of Definition 1.1, and hence uniqueness does not hold. Specifically, one solution is $u(x,t)=u^{*}$ a.e. in $\mathbb{R}^{N}\times[0,\infty)$. The other solution satisfies $u(x,t)\in L^{\infty}_{loc}((0,t_{0}),L^{\infty}(\mathbb{R}^{N}))$ for some $t_{0}>0$, $u(x,t)$ satisfies problem (1.1) in the classical sense for $(x,t)\in\mathbb{R}^{N}\times(0,t_{0})$ and
(1.6) $\lim_{t\to 0^{+}}\left\|u(t)-u^{*}\right\|_{L^{\gamma}_{ul}(\mathbb{R}^{N})}=0,$
provided that $1\leq\gamma<\gamma^{*}$, where $\gamma^{*}$ is defined in Proposition 1.2.
Note that if $q_{f}=1$, then (A5) is equivalent to $2<N<10$.
In the pure power case $f(u)=u^{\beta}$ the conclusion of Theorem 1.4 holds for $p_{S}<\beta<p_{JL}$. The model case in the present paper is $f(u)=u^{\beta}e^{u^{\gamma}}$, $\beta\geq p_{S}$ and $\gamma>1$. In this case we see in Section 2 that the conclusion of Theorem 1.4 holds for $2<N<10$.
The second main theorem is nonuniqueness for critical and subcritical cases.
Theorem 1.5 (Nonuniqueness (critical/subcritical)).
Let $N>2$, and let
$p_{0}:=\frac{N}{N-2}\quad\text{and}\quad q_{0}:=\frac{p_{0}}{p_{0}-1}=\frac{N}{2}.$
Suppose (A1)(A3) and the following (A7) and (A8):
  • (A7)
    $q_{S}\leq q_{f}<q_{0}$ (equivalently, $p_{0}<p_{f}\leq p_{S}$),
  • (A8)
    Problem (1.1) has a positive radial singular stationary solution $u^{*}$ such that
[…]
Then, problem (1.1) with $u_{0}=u^{*}$ has at least two positive solutions in the sense of Definition 1.1, and hence uniqueness does not hold. Specifically, one solution is $u(x,t)=u^{*}$ a.e. in $\mathbb{R}^{N}\times[0,\infty)$. The other solution satisfies $u(x,t)\in L^{\infty}_{loc}((0,t_{0}),L^{\infty}(\mathbb{R}^{N}))$ for some $t_{0}>0$, $u(x,t)$ satisfies problem (1.1) in the classical sense for $(x,t)\in\mathbb{R}^{N}\times(0,t_{0})$ and (1.6) holds provided that $1\leq\gamma<\gamma^{*}$.
Because of (A8), nonuniqueness problem can be reduced to the following:
Problem 1.6.
Let $N>2$. Assume that (A1)(A3) and (A7) hold. Find conditions on $f$ such that problem (1.1) has a positive radial singular stationary solution $u^{*}$ satisfying (1.7).
Singular solutions satisfying (1.7) are recently constructed near the origin for $p_{0}<p_{f}<p_{S}$ in [9]. The authors of [22] showed that every positive radial solution diverging to $\infty$ as $|x|\to 0$ satisfies (1.7) in the subcritical case $p_{0}<p_{f}<p_{S}$. This will be published elsewhere. On the other hand, it is well-known that $\Delta u+u^{p}=0$ has a positive radial singular solution that does not satisfy (1.7) in the critical case $p=p_{S}$ ($q=q_{S}$). Note that in the supercritical case the singular solution is unique and it always satisfies (1.7) (Proposition 3.2). Therefore, the case $p=p_{S}$ is exceptional in the range $p>p_{0}$ ($q<q_{0}$).
Even if a singular solution $u^{*}$ is constructed near the origin, it is not easy to determine whether $u^{*}(r)$ has a zero or not in the case $p_{0}<p_{f}<p_{S}$. Pohozaev’s identity does not seem to work. Problem 1.6 is an open question at this moment.
Let us recall previous research about nonuniqueness. Let $N>2$. The pure power case
(1.8) $\begin{cases}\partial_{t}u-\Delta u=|u|^{p-1}u,&x\in\mathbb{R}^{N},\ t>0,\\ u(x,0)=u_{0}(x),&x\in\mathbb{R}^{N}\end{cases}$
has been intensively studied by many researchers. The equation (1.8) is invariant under the transformation
$u(x,t)\mapsto u_{\lambda}(x,t)=\lambda^{\frac{2}{p-1}}u(\lambda x,\lambda^{2}t)\quad\text{for all}\ \lambda>0.$
We call a solution $u$ of problem (1.8) the forward self-similar solution if $u(x,t)=u_{\lambda}(x,t)$ for all $\lambda>0$. Let $\gamma_{c}$ be defined by
$\gamma_{c}:=\frac{N(p-1)}{2}.$
Note that $\gamma^{*}={N/2(q_{f}-1)}={N(p_{f}-1)/2}=\gamma_{c}$ if $p=p_{f}<\infty$. Haraux-Weissler [14] constructed a positive forward self-similar solution $u(t)$ of problem (1.8) that is in $C([0,T],L^{\gamma}(\mathbb{R}^{N}))$ for $1\leq\gamma<\gamma_{c}$ by finding a positive radial solution of
(1.9) $\Delta\varphi+\frac{1}{2}\xi\cdot\nabla\varphi+\frac{\varphi}{p-1}+|\varphi|^{p-1}\varphi=0.$
Let
$p_{F}:=1+\frac{2}{N}.$
If $p_{F}<p<p_{S}$, then $\left\|u(t)\right\|_{L^{\gamma}(\mathbb{R}^{N})}\to 0$ as $t\to 0^{+}$, and hence problem (1.8) with $u_{0}=0$ has two solutions, namely $u(t)$ and the trivial one. Therefore, uniqueness does not hold in $C([0,T],L^{\gamma}(\mathbb{R}^{N}))$. Since self-similar solutions are closely related to nonuniqueness of solutions and asymptotic behaviors of solutions to problem (1.8), the elliptic equation (1.9) has attracted much attention. The details of the structure of radial solutions of equation (1.9) can be found in [24]. Nonradial nonnegative solutions of equation (1.9) were constructed for $p_{F}<p<p_{S}$ in Naito [23]. If $1<p\leq p_{F}$, then it is well-known that problem (1.8) does not have global-in-time positive solutions (see [10]). Therefore, problem (1.9) has no positive solution for $1<p\leq p_{F}$, since a self-similar solution gives a global-in-time solution. For sign-changing radial solutions of equation (1.9) in the case $p_{F}<p<p_{S}$, see Cazenave et.al. [2].
Let $B\subset\mathbb{R}^{N}$, $N>2$, denote the unit ball. Ni-Sacks [25] studied Cauchy problem
(1.10) $\begin{cases}\partial_{t}u-\Delta u=u^{p_{0}},&x\in B,\ t>0,\\ u=0,&x\in\partial B,\ t>0,\\ u(x,0)=u_{0},&x\in B.\end{cases}$
They constructed a positive singular stationary solution $u^{*}$ that belongs to $L^{p_{0}}(B)$. Moreover, they showed that problem (1.10) with $u_{0}=u^{*}$ has a positive bounded solution $u(t)\in C([0,T],L^{p_{0}}(B))\cap L^{\infty}_{loc}((0,T),L^{\infty}(\mathbb{R}^{N}))$. Therefore, uniqueness of positive solutions in $C([0,T],L^{p_{0}}(B))$ does not hold. When the domain is $\mathbb{R}^{N}$, nonuniqueness of positive solutions of problem (1.8) in $C([0,T],L^{p_{0}}(\mathbb{R}^{N}))$ was proved in Terraneo [31]. Later, infinitely many solutions of problem (1.8) in $C([0,T],L^{p_{0}}(\mathbb{R}^{N}))$ with one initial data were constructed in Matos-Terraneo [19] and Takahashi [29].
Let $u^{*}$ denote the explicit positive radial singular stationary solution of problem (1.8), i.e.,
(1.11) $u^{*}(x)=L|x|^{-\frac{2}{p-1}},\quad L:=\left\{\frac{2}{p-1}\left(N-2-\frac{2}{p-1}\right)\right\}^{\frac{1}{p-1}}$
for $p>p_{0}$. In Galaktionov-Vazquez [11] a positive bounded solution of problem (1.8) with $u_{0}=u^{*}$ in the sense of the proper solution was constructed for $p_{S}<p<p_{JL}$, and hence uniqueness does not hold. The definition and properties of the proper solution can be found in [11, Section 2]. They also studied Cauchy-Dirichlet problem
(1.12) $\begin{cases}\partial_{t}u-\Delta u=2(N-2)e^{u},&x\in B,\ t>0,\\ u=0,&x\in\partial B,\ t>0,\\ u(x,0)=u_{0},&x\in B.\end{cases}$
This problem has the explicit positive singular stationary solution $u^{*}(|x|)=-2\log|x|$. If $2<N<10$, then a bounded solution of problem (1.12) with $u_{0}=u^{*}$ in the sense of the proper solution was constructed in [11]. Hence, uniqueness does not hold for $2<N<10$. See also [26] for problem (1.12).
When the domain is a ball in $\mathbb{R}^{2}$, Ioku et.al. [17] and Ibrahim et.al. [15] proved nonuniqueness of positive solutions of Cauchy-Dirichlet problem. Specifically, they constructed a positive singular stationary solution $u^{*}$ and a bounded solution with initial data $u^{*}$ when the nonlinearity $f$ has an exponential growth.
Nonuniqueness was studied for the equation $\partial_{t}u=\Delta u+|x|^{\beta}|u|^{p-1}u$ in [30, 3, 12]. For nonuniqueness for other equations, see [15, Remark 3] and references therein.
Theorem 1.4 unifies two nonuniqueness results of [11], since $p_{S}<p_{f}<p_{JL}$ in the power case ($p_{f}\neq 1$) and $2<N<10$ in the super-power case $(p_{f}=1)$. However, we will construct the solution defined in Definition 1.1 and do not directly use the proper solution.
Let us recall previous research about uniqueness. A natural question is to find the largest function spaces such that uniqueness holds. Uniqueness of solutions of problem (1.8) have been studied by many researchers. Brezis-Cazenave [1] showed that the solution of problem (1.8), which was constructed in Weissler [32], is unique in $C([0,T],L^{\gamma}(\mathbb{R}^{N}))$ if either $\gamma=\gamma_{c}>p$ or both $\gamma>\gamma_{c}$ and $\gamma\geq p$. When $\gamma$ is in these ranges, we see that $u^{*}(x)\chi_{B}(x)\not\in L^{\gamma}(\mathbb{R}^{N})$ and hence the singularity of $u^{*}$ is too strong to obtain uniqueness. As mentioned above, uniqueness in $C([0,T],L^{p_{0}}(\mathbb{R}^{N}))$ does not hold for $p=p_{0}$. We remark that $p=p_{0}\iff\gamma_{c}=p$, and hence this case is a borderline case of [1].
We briefly introduce the Lorentz space $L^{p,\infty}(\mathbb{R}^{N})$ to mention the uniqueness result of Chikami et.al. [3]. For $0<p\leq\infty$,
$L^{p,\infty}(\mathbb{R}^{N}):=\left\{u\in M(\mathbb{R}^{N});\ \left\|u\right\|_{L^{p,\infty}(\mathbb{R}^{N})}:=\sup_{\lambda>0}\lambda\left|\left\{x\in\mathbb{R}^{N};\ |u(x)|>\lambda\right\}\right|^{\frac{1}{p}}<\infty\right\},$
where $M(\mathbb{R}^{N})$ denote the set of measurable functions on $\mathbb{R}^{N}$ and $|S|$ denotes the Lebesgue measure of the set $S\subset\mathbb{R}^{N}$. The space $\mathcal{L}^{p,\infty}(\mathbb{R}^{N})$ is the closure of $L^{p,\infty}(\mathbb{R}^{N})\cap L^{\infty}(\mathbb{R}^{N})$ in the space $L^{p,\infty}(\mathbb{R}^{N})$. The uniqueness of the solution of (1.8) in $C([0,T],\mathcal{L}^{\gamma_{c},\infty}(\mathbb{R}^{N}))$ was proved for $p>p_{S}$ in [3, Theorem 1.3 i]. We can check that $u^{*}\in L^{\gamma_{c},\infty}(\mathbb{R}^{N})\setminus\mathcal{L}^{\gamma_{c},\infty}(\mathbb{R}^{N})$, and hence their theorem cannot apply to the case $u_{0}=u^{*}$.
In the case of the general nonlinearity (1.1) there are few theorems about uniqueness, although existence and nonexistence theorems were recently developed in [8, 6, 7].
The case $N=2$ is out of our scope. However, in Ioku et.al. [16] the uniqueness of the solution of problem (1.1) in $C([0,T],{\rm exp}L_{0}^{2}(\mathbb{R}^{2}))$ was established if $f$ satisfies
$|f(x)-f(y)|\leq C|x-y|\left(e^{\lambda x^{2}}+e^{\lambda y^{2}}\right).$
Here, ${\rm exp}L_{0}^{2}(\mathbb{R}^{2})$ is the closure of $C_{0}^{\infty}(\mathbb{R}^{2})$ in the Orlicz space ${\rm exp}L^{2}(\mathbb{R}^{2})$. The uniqueness of the positive radial singular solution on $B$ is an open problem. It was proved in [16, 17] that for a certain positive singular stationary solution $u^{*}$ on $B$ of a certain explicit nonlinearity $f$, $u^{*}\in{\rm exp}L^{2}(B)\setminus{\rm exp}L^{2}_{0}(B)$.
Remark 1.7.
As mentioned above, Fujishima et.al. [7] obtained rather sharp existence and nonexistence criteria. It follows from [7, Corollary 1.1] that there exists a unique $\lambda_{0}>0$ such that problem (1.8) with $u_{0}(x)=F^{-1}[{\lambda^{-1}|x|^{2}}]\chi_{B}(x)$ has a solution for $0<\lambda<\lambda_{0}$ and no solution for $\lambda>\lambda_{0}$. However, the exact value of $\lambda_{0}$ is an open problem. Theorem 1.4 indicates that $\lambda_{0}\geq 2N-4q_{f}$ if $q_{JL}<q<q_{S}$. When $q_{f}=1$, the authors of [13] proved that $\lambda_{0}=2N-4$ under additional assumptions on solutions and the nonlinearity. When $f(u)=u^{p}$, $p\geq p_{JL}$, it was proved in [11] that $\lambda_{0}=2N-4q$ in the framework of the proper solution, where q:=p/(p1)q:=p/(p-1). If $p_{S}<p<p_{JL}$, then $\lambda_{0}>2N-4q$ by Souplet-Weissler [28].
Let us mention technical details. We will see in Section 3 that the positive singular solution $u^{*}$ given in Proposition 1.2 is a positive singular stationary solution of problem (1.1) in the sense of Definition 1.1. We also construct a positive bounded solution $u(t)$ of problem (1.1) with $u_{0}=u^{*}$. This solution is in $L^{\infty}_{loc}((0,T),L^{\infty}(\mathbb{R}^{N}))$, and hence $u(t)$ is different from $u^{*}$.
The method of the construction of $u(t)$ is based on the comparison principle and not on the contraction mapping theorem. We construct a supersolution $v(t)$ defined by (4.17) with (4.6), combining the singular solution and a transformation of a self-similar solution. Let $u_{0n}:=\min\{u^{*},n\}$. Then, $u_{0n}\in BUC(\mathbb{R}^{N})$, and hence (1.1) with $u_{0}=u_{0n}$ has a classical solution $u_{n}(t)$ for $0\leq t<t_{0}$. Since $u_{n}$ is a classical solution, the maximum principle works and we obtain the monotone sequence
$u_{1}(t)<u_{2}(t)<\cdots<u_{n}(t)<\cdots<v(t)$
for $0<t<t_{0}$. Here, the inequality $u_{n}(t)<v(t)$ follows from the definition of $u_{n}$. Letting $n\to\infty$, we obtain the limit function
$u(t):=\lim_{n\to\infty}u_{n}(t),$
which satisfies (1.4) by the monotone convergence theorem. Since $u(t)\leq v(t)$ and $v(t)$ is bounded for each $t>0$, $u(t)$ is the desired bounded solution. This method is inspired by the definition of the proper solution which is explained in [11]. However, we have to construct a bounded solution $u(t)$ in the sense of Definition 1.1.
This paper consists of six sections. In Section 2 we give three examples of $f$. In Section 3 we construct a positive radial singular solution $u^{*}$ and prove Proposition 1.2 and Lemma 1.3. In Section 4 we recall basic properties of forward self-similar solutions of problem (1.1) in the cases $f(u)=u^{p}$ and $e^{u}$. Combining the self-similar solution and the singular stationary solution, we construct a supersolution $v$ of problem (1.1) in the sense of Definition 1.1. In Section 5 we construct a bounded solution of problem (1.1) with $u_{0}=u^{*}$, using the supersolution $v$. Then we prove Theorem 1.4. In Section 6 we briefly prove Theorem 1.5, modifying the proof of Theorem 1.4.
2. Examples
2.1. Example 1
The first example is
$f(u)=u^{\beta}e^{u^{\gamma}},\quad\beta\geq p_{S},\quad\gamma>1.$
We easily see that (A1), (A2) and (A4) hold. See [13, Example 1] for details. By direct calculation we have
(2.1) $\frac{f^{\prime}(u)^{2}}{f(u)f^{\prime\prime}(u)}=\frac{\beta^{2}+2\beta\gamma u^{\gamma}+\gamma^{2}u^{2\gamma}}{\beta(\beta-1)+\gamma(2\beta+\gamma-1)u^{\gamma}+\gamma^{2}u^{2\gamma}},$
and hence (A3) with $q_{f}=1$ holds. Moreover, the RHS of (2.1) is less than or equal to $1$ for large $u>0$, and hence (A6) holds. If $2<N<10$, then $q_{JL}<q_{f}=1<q_{S}$, and hence (A5) holds. Therefore, the conclusion of Theorem 1.4 holds for $2<N<10$.
2.2. Example 2
The second example is
$f(u)=\chi(u)e^{au},\quad a=20,$
where $\chi(u):=\int_{0}^{u}\chi^{\prime}(s)ds$ and
$\chi^{\prime}(u):=\begin{cases}0&\text{for}\ u\geq 4;\\ 5(u-4)^{4}&\text{for}\ 3\leq u\leq 4;\\ 10-5(u-2)^{4}&\text{for}\ 1\leq u\leq 3;\\ 5u^{4}&\text{for}\ 0\leq u\leq 1.\end{cases}$
In particular, $f\in C^{1}[0,\infty)\cap C^{2}(0,\infty)$ and
$f(u)=\begin{cases}u^{5}e^{au}&\text{for}\ 0\leq u\leq 1;\\ 20e^{au}&\text{for}\ u\geq 4.\end{cases}$
We easily see that (A1), (A2) and (A4) hold. See [13, Example 2] for details. By direct calculation we have
$\frac{f^{\prime}(u)^{2}}{f(u)f^{\prime\prime}(u)}=1,$
for large $u>0$, and hence (A3) with $q_{f}=1$ and (A6) hold. If $2<N<10$, then $q_{JL}<q_{f}=1<q_{S}$, and hence (A5) holds. Therefore, the conclusion of Theorem 1.4 holds for $2<N<10$.
2.3. Example 3
The third example is
$f(u)=u^{\beta}+u^{\gamma},\quad p_{S}<\gamma<\beta<p_{JL}.$
It is easy to check that (A1), (A2) and (A4) hold. Since
$\frac{f^{\prime}(u)^{2}}{f(u)f^{\prime\prime}(u)}=\frac{\beta^{2}u^{2(\beta-\gamma)}+2\beta\gamma u^{\beta-\gamma}+\gamma^{2}}{\beta(\beta-1)u^{2(\beta-\gamma)}+(\beta^{2}-\beta+\gamma^{2}-\gamma)u^{\beta-\gamma}+\gamma(\gamma-1)},$
(A3) with $q_{f}=\beta/(\beta-1)$ holds, and hence $p_{f}=\beta$. Since $p_{S}<p_{f}<p_{JL}$, we see that $q_{JL}<q_{f}<q_{S}$, and hence (A5) holds. Since $q_{f}\neq 1$, (A6) automatically holds. Therefore, the conclusion of Theorem 1.4 holds.
3. Singular solution
Lemma 3.1.
Suppose that (A1)(A3). Then,
$\lim_{u\to\infty}f^{\prime}(u)F(u)=q_{f}.$
Proof.
It is obvious that
$F(u)\to 0\quad\text{as}\ u\to\infty.$
By (A1) and (A2) we see that $f^{\prime}(u)\to\infty$ as $u\to\infty$. Hence
$\frac{1}{f^{\prime}(u)}\to 0\quad\text{as}\ u\to\infty.$
By L’Hospital’s rule we have
$\lim_{u\to\infty}\frac{F(u)}{1/f^{\prime}(u)}=\lim_{u\to\infty}\frac{-1/f(u)}{-f^{\prime\prime}(u)/f^{\prime}(u)^{2}}=\lim_{u\to\infty}\frac{f^{\prime}(u)^{2}}{f(u)f^{\prime\prime}(u)}=q_{f}.$
Hereafter in this section we consider the equation
(3.1) $u^{\prime\prime}+\frac{N-1}{r}u^{\prime}+f(u)=0.$
We call a solution $u^{*}$ of equation (3.1) a singular solution if $u^{*}(r)\to\infty$ as $r\to 0^{+}$.
Let $u(r,\alpha)$, $\alpha>0$, denote a solution of the initial value problem (3.1) with $u(0,\alpha)=\alpha$ and $u_{r}(0,\alpha)=0$. Because of Lemma 3.1 and (A1)–(A3), all assumptions of [21, Theorem 1.1 and Lemma 2.5], which are stated in Proposition 3.2, are satisfied.
Proposition 3.2.
Suppose that $N>2$ and that (A1)(A3) hold. If $q_{f}<q_{S}$, then there exists a unique positive singular solution $u^{*}$ of equation (3.1) for $0<r\leq r_{1}$ with some $r_{1}>0$, and the regular solution $u(r,\alpha)$ satisfies
$u(r,\alpha)\to u^{*}(r)\ \ \textrm{in}\ \ C^{2}_{loc}(0,r_{1}]\ \ \textrm{as}\ \ \alpha\to\infty.$
Furthermore, the positive singular solution $u^{*}$ satisfies
(3.2) $r^{N-1}u^{*\prime}(r)=\int_{0}^{r}f(u^{*}(s))s^{N-1}\,\mathop{}\!\mathrm{d}s$
for $0<r<r_{1}$ and
(3.3) $u^{*}(r)=F^{-1}\left[\frac{r^{2}}{2N-4q_{f}}(1+\theta(r))\right]\ \ \textrm{as}\ \ r\to 0^{+},$
where $F^{-1}$ is the inverse function of $F(u)$ defined by (1.2) and $\theta(r)$ satisfies $\lim_{r\to 0^{+}}\theta(r)=0$,
(3.4) $\theta(r)\in C^{1}(0,r_{1}),\quad\text{and}\quad\lim_{r\to 0^{+}}r\theta^{\prime}(r)=0.$
All statements of Proposition 3.2 except (3.4) are the same as [21, Theorem 1.1 and Lemma 2.5]. We briefly prove (3.4). Because of the uniqueness of the singular solution, $u^{*}$ given in Proposition 3.2 and $u^{*}$ constructed in [20, Theorem A] are identical. Let $t:=\log r$ and $x(t):=\theta(r)=-1+(2N-4q_{f})r^{-2}F[u^{*}(r)]$. The solution constructed in [20, Lemma 3.1] satisfies $x(t)$ is of class $C^{1}$ and $x^{\prime}(t)\to 0$ as $t\to$-$\infty$. Then, we obtain
$r\theta^{\prime}(r)=x^{\prime}(t)\to 0\quad\text{as}\ t\to-\infty.$
Thus, (3.4) holds.
In the authors’ previous paper [13] a singular stationary solution was constructed in the case $q_{f}=1$. Proofs of Lemmas 3.3, Proposition 1.2 and Lemma 1.3 in the present paper, which treat the case $q_{f}<q_{S}$, are slight modifications of those of [13, Lemma 3.3, Proposition 1.2 and Lemma 4.2], respectively. However, we show proofs for self-contained purpose.
Lemma 3.3.
Suppose that $N>2$ and that (A1)(A3) hold. Let $u^{*}$ be the unique positive singular solution of equation (3.1) on $0<r<r_{1}$ given in Proposition 3.2. For any small $\delta>0$, there exist $R_{0}\in(0,r_{1})$ and $C>0$ such that
(3.5) $u^{*}(r)\leq Cr^{2-2q_{f}-2\delta}\quad\mbox{and}\quad|u^{*\prime}(r)|\leq Cr^{1-2q_{f}-2\delta}$
for $r\in(0,R_{0})$.
Proof.
Let $\delta>0$ be sufficiently small. By Lemma 3.1 we have
$\lim_{u\to\infty}f^{\prime}(u)F(u)=q_{f}.$
There exists $M>0$ such that $f^{\prime}(u)F(u)\leq q_{f}+\delta$ for all $u\geq M$. Then
$\frac{f^{\prime}(u)}{f(u)}\leq\frac{q_{f}+\delta}{f(u)F(u)}=-\frac{(q_{f}+\delta)F^{\prime}(u)}{F(u)}$
for all $u\geq M$. Integrating this inequality in $u$ over $(M,u)$, we have
$\log\frac{f(u)}{f(M)}\leq-(q_{f}+\delta)\log\frac{F(u)}{F(M)}$
for all $u\geq M$. Therefore,
(3.6) $\frac{1}{f(M)F(M)^{q_{f}+\delta}}\leq\frac{1}{f(u)F(u)^{q_{f}+\delta}}=-\frac{F^{\prime}(u)}{F(u)^{q_{f}+\delta}}$
for all $u\geq M$. Since $u^{*}(r)\to\infty$ as $r\to 0^{+}$, there exists $r_{2}>0$ such that
$u^{*}(r)>M\quad\mbox{for}\ r\in(0,r_{2}).$
Integrating (3.6) in $u$ over $(M,u^{*}(r))$, we have
$u^{*}(r)<\frac{f(M)F(M)^{q_{f}+\delta}}{q_{f}-1+\delta}(F(u^{*}(r))^{1-q_{f}-\delta}-F(M)^{1-q_{f}-\delta})+M$
for $r\in(0,r_{2})$. Since $F(u^{*}(r))=(2N-4q_{f})^{-1}r^{2}(1+o(1))$ as $r\to 0^{+}$ (by (3.3)), there exist $R_{0}\in(0,r_{2})$ and $C>0$ such that
$u^{*}(r)\leq Cr^{2-2q_{f}-2\delta}\quad\mbox{for}\ r\in(0,R_{0}).$
By (3.6) we have
(3.7) $f(u^{*}(r))<\frac{f(M)F(M)^{q_{f}+\delta}}{F(u^{*}(r))^{q_{f}+\delta}}\leq Cr^{-2q_{f}-2\delta}$
for $r\in(0,R_{0})$. By (3.2) we have
$0<-r^{N-1}u^{*\prime}(r)=\int_{0}^{r}f(u^{*}(s))s^{N-1}\,\mathop{}\!\mathrm{d}s\leq C\int_{0}^{r}s^{N-1-2q_{f}-2\delta}\,\mathop{}\!\mathrm{d}s\leq Cr^{N-2q_{f}-2\delta}.$
Then we have
$|u^{*\prime}(r)|\leq Cr^{1-2q_{f}-2\delta}\quad\mbox{for}\quad r\in(0,R_{0}).$
The proof is complete. ∎
Proof of Proposition 1.2.
Let $u^{*}$ denote the unique positive singular solution of equation (3.1) on $0<r<r_{1}$ given in Proposition 3.2. We extend the domain of $u^{*}$ such that $u^{*}$ satisfies equation (3.1) for all $r>0$.
We show that $u^{*}(r)>0$ for $r>0$. Let $P(u^{*}(r))$ be defined by
$P(u^{*}(r)):=\frac{1}{2}r^{N}u^{*\prime}(r)^{2}+r^{N}F_{0}(u^{*}(r))+\frac{N-2}{2}r^{N-1}u^{*}(r)u^{*\prime}(r)$
for $r>0$, where
$F_{0}(u):=\int_{0}^{u}f(s)\,\mathop{}\!\mathrm{d}s.$
Using (3.1), we have
$\frac{d}{dr}P(u^{*}(r))=-\frac{N-2}{2}r^{N-1}Q(u^{*}(r)),$
where $Q$ is defined in (A4). Integrating the above equality over $(\rho,r)$, we have
(3.8) $P(u^{*}(r))-P(u^{*}(\rho))=-\frac{N-2}{2}\int_{\rho}^{r}s^{N-1}Q(u^{*}(s))\,\mathop{}\!\mathrm{d}s.$
Since $q_{f}<q_{S}$, we can take a small $\delta>0$ such that $N+2-4q_{f}-4\delta>0$. By L’Hospital’s rule we have
$\lim_{\rho\to 0^{+}}\left|\frac{F_{0}(u^{*}(\rho))}{\rho^{-N}}\right|=\lim_{\rho\to 0^{+}}\left|\frac{f(u^{*}(\rho))u^{*\prime}(\rho)}{-N\rho^{-N-1}}\right|\leq\lim_{\rho\to 0^{+}}C\rho^{N+2-4q_{f}-4\delta}=0,$
where Lemma 3.3 and (3.7) are used. By the definition of $P$ we have
$|P(u^{*}(\rho))|$ $\leq C\rho^{N+2-4q_{f}-4\delta}+|\rho^{N}F_{0}(u^{*}(\rho))|+C\rho^{N+2-4q_{f}-4\delta}$
$\to 0\quad\mbox{as}\quad\rho\to 0^{+}.$
Assume that there exists $r_{0}>0$ such that $u^{*}(r_{0})=0$. Then, by Hopf’s lemma, $u^{*\prime}(r_{0})<0$. Letting $\rho\to 0^{+}$ in (3.8) with $r=r_{0}$, we have
$0<P(u^{*}(r_{0}))-0=-\frac{N-2}{2}\int_{0}^{r_{0}}s^{N-1}Q(u^{*}(s))\,\mathop{}\!\mathrm{d}s\leq 0,$
which is a contradiction. Thus, $r_{0}=\infty$, and hence $u^{*}(r)>0$ for $0\leq r<\infty$.
Next, assume that there exists $r_{0}^{\prime}>0$ such that $u^{*\prime}(r_{0}^{\prime})=0$. Since $u^{*}(r_{0}^{\prime})>0$, we have
$P(u^{*}(r_{0}^{\prime}))=r_{0}^{\prime N}F_{0}(u^{*}(r_{0}^{\prime}))>0.$
Therefore, we arrive at the same conclusion as the above inequality with $r_{0}$ replaced by $r_{0}^{\prime}$, which is a contradiction. Thus, $u^{*\prime}$ does not change its sign on $(0,\infty)$, and hence $u^{*\prime}(r)<0$ for all $r>0$ since $u^{*\prime}(r)\to\infty$ as $r\to 0^{+}$. We can easily see that the other properties in (1.5) are satisfied.
Hereafter, we show that $u^{*}\in\mathcal{L}^{\gamma}_{{ul}}(\mathbb{R}^{N})$ for $1\leq\gamma<\gamma^{*}$. Since $u^{*\prime}(r)<0$ for all $r>0$, at least $\limsup_{r\to\infty}|u^{*}(r)|<\infty$ holds. Thus, it suffices to consider integrability only near the origin. Let $u_{n}(x):=\min\{n,u^{*}(x)\}$ for $n\geq 1$. Then it is obvious that $u_{n}\in BUC(\mathbb{R}^{N})$. Fix $1\leq\gamma<\gamma^{*}$ arbitrarily. Since $N+\gamma(2-2q_{f})>0$, we can take a sufficiently small $\delta>0$ such that $N+\gamma(2-2q_{f}-2\delta)>0$. By Lemma 3.3 we have
$\int_{B(0,2)}|u^{*}(x)-u_{n}(x)|^{\gamma}\,\mathop{}\!\mathrm{d}x\leq C\int_{0}^{2}r^{N+\gamma(2-2q_{f}-2\delta)-1}\,\mathop{}\!\mathrm{d}r<\infty.$
By the dominated convergence theorem we have
$\lim_{n\to\infty}\int_{B(0,2)}|u^{*}(x)-u_{n}(x)|^{\gamma}\,\mathop{}\!\mathrm{d}x=0,$
which indicates that $\left\|u^{*}-u_{n}\right\|_{L^{\gamma}_{{ul}}(\mathbb{R}^{N})}\to 0$ as $n\to\infty$. Thus, $u^{*}\in\mathcal{L}^{\gamma}_{{ul}}(\mathbb{R}^{N})$ for $1\leq\gamma<\gamma^{*}$. The proof is complete. ∎
We use the following proposition to prove Lemma 1.3.
Proposition 3.4 ([18, Proposition 2.2]).
For $1\leq p<\infty$, let $\mathcal{L}^{p}_{ul}(\mathbb{R}^{N})$ be defined by (1.3). The following assertions are equivalent:
  • (i)
    $u\in\mathcal{L}^{p}_{ul}(\mathbb{R}^{N})$;
  • (ii)
    $\lim_{|y|\to 0^{+}}\left\|u(\,\cdot\,+y)-u(\,\cdot\,)\right\|_{L^{p}_{ul}(\mathbb{R}^{N})}=0$;
  • (iii)
    $\lim_{t\to 0^{+}}\left\|S(t)u-u\right\|_{L^{p}_{ul}(\mathbb{R}^{N})}=0$.
Proof of Lemma 1.3.
Let $x\in\mathbb{R}^{N}\setminus\{0\}$ and $t\in(0,\infty)$. We show that $u(x,t)=u^{*}(x)$ is a solution in the sense of Definition 1.1. Since $u^{*}(x)\in C^{2}(\mathbb{R}^{N}\setminus\{0\})$, we see by Green’s identity that, for $s\in(0,t)$,
(3.9) $\begin{split}&\int_{B(0,\varepsilon)^{c}}G_{t-s}(x-y)\Delta u^{*}(y)\,\mathop{}\!\mathrm{d}y\\ &=\int_{\partial B(0,\varepsilon)^{c}}\left(G_{t-s}(x-y)\frac{\partial}{\partial\nu_{y}}u^{*}(y)-u^{*}(y)\frac{\partial}{\partial\nu_{y}}G_{t-s}(x-y)\right)\,\mathop{}\!\mathrm{d}\sigma(y)\\ &\qquad+\int_{B(0,\varepsilon)^{c}}\Delta_{y}G_{t-s}(x-y)u^{*}(y)\,\mathop{}\!\mathrm{d}y\\ &=:I_{1}+I_{2}.\end{split}$
By Lemma 3.3 we have
$|I_{1}|$ $\leq C\left|\left.\frac{\partial}{\partial\nu}u^{*}(y)\right|_{|y|=\varepsilon}\right|\varepsilon^{N-1}+C\left|\left.u^{*}(y)\right|_{|y|=\varepsilon}\right|\varepsilon^{N-1}$
(3.10) $\leq C\varepsilon^{N-2q_{f}-2\delta}+C\varepsilon^{N+1-2q_{f}-2\delta}\to 0\ \ \textrm{as}\ \ \varepsilon\to 0^{+},$
since $N-2q_{f}-2\delta>0$ and $N+1-2q_{f}-2\delta>0$ for small $\delta>0$. Integrating (3.9) with respect to $s$ over $(0,t-\delta)$ , we have
$\begin{split}&\int_{0}^{t-\delta}\int_{B(0,\varepsilon)^{c}}G_{t-s}(x-y)\Delta u^{*}(y)\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\\ &\qquad=\int_{0}^{t-\delta}I_{1}\,\mathop{}\!\mathrm{d}s-\int_{0}^{t-\delta}\int_{B(0,\varepsilon)^{c}}\partial_{s}G_{t-s}(x-y)u^{*}(y)\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\\ &\qquad=\int_{0}^{t-\delta}I_{1}\,\mathop{}\!\mathrm{d}s-\int_{B(0,\varepsilon)^{c}}\int_{0}^{t-\delta}\partial_{s}G_{t-s}(x-y)u^{*}(y)\,\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}y\\ &\qquad=\int_{0}^{t-\delta}I_{1}\,\mathop{}\!\mathrm{d}s-\int_{B(0,\varepsilon)^{c}}\left(G_{\delta}(x-y)u^{*}(y)-G_{t}(x-y)u^{*}(y)\right)\,\mathop{}\!\mathrm{d}y,\end{split}$
where we used $\Delta_{y}G_{t-s}(x-y)=-\partial_{s}G_{t-s}(x-y)$. Letting $\varepsilon\to 0^{+}$, by the dominated convergence theorem and (3) we have
(3.11) $\lim_{\varepsilon\to 0^{+}}\int_{0}^{t-\delta}\int_{B(0,\varepsilon)^{c}}G_{t-s}(x-y)\Delta u^{*}(y)\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s=-S(\delta)u^{*}+S(t)u^{*}.$
By the monotone convergence theorem we have
(3.12) $\begin{split}&\lim_{\varepsilon\to 0^{+}}\int_{0}^{t-\delta}\int_{B(0,\varepsilon)^{c}}G_{t-s}(x-y)f(u^{*}(y))\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\\ &\qquad=\lim_{\varepsilon\to 0^{+}}\int_{0}^{t-\delta}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)f(u^{*}(y))\chi_{B(0,\varepsilon)^{c}}(y)\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\\ &\qquad=\int_{0}^{t-\delta}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)f(u^{*}(y))\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s.\end{split}$
The solution $u^{*}$ satisfies $\Delta u^{*}+f(u^{*})=0$ in the classical sense for $x\in B(0,\varepsilon)^{c}$. Hence,
$\begin{split}&\int_{0}^{t-\delta}\int_{B(0,\varepsilon)^{c}}G_{t-s}(x-y)\Delta u(y)\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\\ &\qquad\qquad+\int_{0}^{t-\delta}\int_{B(0,\varepsilon)^{c}}G_{t-s}(x-y)f(u^{*}(y))\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s=0.\end{split}$
Letting $\varepsilon\to 0^{+}$, by (3.11) and (3.12) we have
(3.13) $S(\delta)u^{*}+S(t)u^{*}+\int_{0}^{t-\delta}S(t-s)f(u^{*})\,\mathop{}\!\mathrm{d}s=0.$
Since $u^{*}\in\mathcal{L}^{\gamma}_{{ul}}(\mathbb{R}^{N})$ (by Proposition 1.2), by Proposition 3.4 with $p=\gamma$ we have
(3.14) $\lim_{\delta\to 0^{+}}\left\|S(\delta)u^{*}-u^{*}\right\|_{L^{\gamma}_{{ul}}(\mathbb{R}^{N})}=0.$
By the monotone convergence theorem we have
(3.15) $\begin{split}\lim_{\delta\to 0^{+}}\int_{0}^{t-\delta}S(t-s)f(u^{*})\,\mathop{}\!\mathrm{d}s&=\lim_{\delta\to 0^{+}}\int_{0}^{t}\left(S(t-s)f(u^{*})\right)\chi_{(0,t-\delta)}(s)\,\mathop{}\!\mathrm{d}s\\ &=\int_{0}^{t}S(t-s)f(u^{*})\,\mathop{}\!\mathrm{d}s.\end{split}$
By (3.15), (3.14), and (3.13) we see that as $\delta\to 0^{+}$,
$u^{*}=S(t)u^{*}+\int_{0}^{t}S(t-s)f(u^{*})\,\mathop{}\!\mathrm{d}s$
for all $(x,t)\in(\mathbb{R}^{N}\setminus\{0\})\times(0,\infty)$. Thus, $u^{*}$ is a solution in the sense of Definition 1.1. The proof is complete. ∎
4. Transformations of self-similar solutions
In order to treat two canonical nonlinearities in a unified way we define
$f_{q}(U):=\begin{cases}U^{p}&\text{if}\ q>1,\\ e^{U}&\text{if}\ q=1,\end{cases}\qquad F_{q}(U):=\int_{U}^{\infty}\frac{\mathop{}\!\mathrm{d}s}{f_{q}(s)}=\begin{cases}\displaystyle{\frac{1}{p-1}U^{-p+1}}&\text{if}\ q>1,\\ e^{-U}&\text{if}\ q=1,\end{cases}$
where p:=q/(q1)p:=q/(q-1). We consider the canonical equation
(4.1) $\partial_{t}U-\Delta U=f_{q}(U).$
For $\lambda>0$, $U_{\lambda}(x,t)$ is defined by the relation
(4.2) $F_{q}[U_{\lambda}(x,t)]=\lambda^{-2}F_{q}[U(\lambda x,\lambda^{2}t)].$
If $U(x,t)$ satisfies (4.1), then $U_{\lambda}$ also satisfies (4.1). We can easily check that
$U_{\lambda}(x,t)=\begin{cases}\lambda^{\frac{2}{p-1}}U(\lambda x,\lambda^{2}t)&\text{if}\ q>1,\\ U(\lambda x,\lambda^{2}t)-2\log\lambda&\text{if}\ q=1.\end{cases}$
To derive a forward self-similar solution we put $\lambda=1/\sqrt{t}$ into (4.2). We redefine $U$ by
(4.3) $F_{q}[U(x,t)]=tF_{q}\left[\varphi(\xi)\right],\quad\xi:=\frac{x}{\sqrt{t}}.$
Since $U$ satisfies (4.1), $\varphi(\xi)$ satisfies
(4.4) $\Delta\varphi+\frac{1}{2}\xi\cdot\nabla\varphi+f_{q}(\varphi)F_{q}(\varphi)+f_{q}(\varphi)=0.$
Specifically, $\varphi$ satisfies
(4.5) $\begin{cases}\displaystyle{\Delta\varphi+\frac{1}{2}\xi\cdot\nabla\varphi+\frac{\varphi}{p-1}+\varphi^{p}=0}&\text{if}\ q>1,\\ \displaystyle{\Delta\varphi+\frac{1}{2}\xi\cdot\nabla\varphi+1+e^{\varphi}=0}&\text{if}\ q=1.\end{cases}$
In summary, if $\varphi$ is a solution of equation (4.4), then $U$ is a self-similar solution of equation (4.1).
Self-similar solutions are important in the study of equation (4.1), e.g., existence of global-in-time solutions, asymptotic behaviors and nonuniqueness. Then, various properties of solutions of equation (4.5) have been studied in numerous papers. We recall known results about the intersection number of radial regular and singular solutions of equation (4.5). Here, we state known results in a unified way, using (4.4). Equation (4.1) has an explicit singular solution
$U^{*}(x)=F_{q}^{-1}\left[\frac{|x|^{2}}{2N-4q}\right]=\begin{cases}L|x|^{-\frac{2}{p-1}}&\text{if}\ q>1,\\ -2\log|x|+\log 2(N-2)&\text{if}\ q=1,\end{cases}$
where $L$ is defined by (1.11) and $F_{q}^{-1}$ denotes the inverse function of $F_{q}$, i.e.,
$F_{q}^{-1}(u)=\begin{cases}\displaystyle{\frac{1}{\{(p-1)u\}^{1/(p-1)}}}&\text{if}\ q>1,\\ -\log u&\text{if}\ q=1.\end{cases}$
The function
$\varphi^{*}(\xi):=F_{q}^{-1}\left[\frac{|\xi|^{2}}{2N-4q}\right],$
which is the same function as $U^{*}$, is a solution of equation (4.4). Let $\eta:=|\xi|$, and let $\varphi(\eta,\alpha)$ denote the solution of
$\begin{cases}\displaystyle{\varphi^{\prime\prime}+\frac{N-1}{\eta}\varphi^{\prime}+\frac{1}{2}\eta\varphi^{\prime}+f_{q}(\varphi)F_{q}(\varphi)+f_{q}(\varphi)=0},&\text{for}\ \eta>0,\\ \varphi(0,\alpha)=\alpha>0,\ \varphi_{\eta}(0,\alpha)=0.\end{cases}$
Proposition 4.1.
Assume that $1\leq q<q_{S}$. Then $\varphi(\eta,\alpha)$ is defined for $\eta\geq 0$ and $\varphi_{\eta}(\eta,\alpha)<0$ for $\eta>0$. There exists $\alpha_{0}>0$ such that $\varphi(\eta,\alpha_{0})$ and $\varphi^{*}(\eta)$ has the first intersection point $\eta_{0}$. Specifically, there exists $\delta>0$ such that
$\begin{cases}\varphi(\eta,\alpha_{0})<\varphi^{*}(\eta)&\text{for}\ 0<\eta<\eta_{0},\\ \varphi(\eta_{0},\alpha_{0})=\varphi^{*}(\eta_{0})\ \text{and}\ \varphi_{\eta}(\eta_{0},\alpha_{0})>\varphi^{*\prime}(\eta_{0}),\\ \varphi(\eta,\alpha_{0})>\varphi^{*}(\eta)&\text{for}\ \eta_{0}<\eta<\eta_{0}+\delta.\end{cases}$
The proof can be found in [28, Theorem 1.4] for $1<q<q_{S}$ and in [5, Proposition 4.2] for $q=1$. The detailed solution structure of the case $1<q<q_{S}$ can be found in [24].
Hereafter, we consider problem (1.1), and let $q_{f}$ be the exponent defined in (A3). If $q_{f}=1$, then we define $q:=1$. If $q_{f}>1$, then we define $q:=q_{f}+\varepsilon$, where $\varepsilon>0$ is determined later. We define $\bar{u}(x,t)$ by
(4.6) $\bar{u}(x,t):=F^{-1}[F_{q}[U(x,t)]]=F^{-1}\left[tF_{q}\left[\varphi\left(\frac{|x|}{\sqrt{t}}\right)\right]\right],$
where $U$ is a self-similar solution of equation (4.1) defined by (4.3), $\varphi(\eta):=\varphi(\eta,\alpha_{0})$ and $\alpha_{0}$ is given in Proposition 4.1. It is obvious from the definition that $\|\overline{u}(\cdot,t)\|_{L^{\infty}(\mathbb{R}^{N})}<\infty$ for every $t>0$. Since
(4.7) $F[\bar{u}(x,t)]=F_{q}[U(x,t)],$
$\bar{u}$ is a transformation of a self-similar solution $U$ of (4.1). Since $U$ satisfies (4.1), by (4.7) we see that $\bar{u}$ is a solution of the equation
(4.8) $\partial_{t}\bar{u}-\Delta\bar{u}=f(\bar{u})+\frac{q-f^{\prime}(\bar{u})F(\bar{u})}{f(\bar{u})F(\bar{u})}|\nabla\bar{u}|^{2}.$
The transformation (4.7) and the equation (4.8) were found in Fujishima-Ioku [8]. If $q_{f}=1$, then by (A6) we integrate $1/f(\bar{u})\leq f^{\prime\prime}(\bar{u})/f^{\prime}(\bar{u})^{2}$ over $\bar{u},\infty)$. Then, $F(\bar{u})\leq/f^{\prime}(\bar{u})$ for large $\bar{u}>0$. Hence, $f^{\prime}(\bar{u})F(\bar{u})\leq q(=1)$ for large $\bar{u}>0$. If $q_{f}>1$, then by Lemma 3.1 we see that $f^{\prime}(\bar{u})F(\bar{u})\to q_{f}$ $(\bar{u}\to\infty)$, and hence $f^{\prime}(\bar{u})F(\bar{u})\leq q(=q_{f}+\varepsilon)$ for large $\bar{u}>0$. In both cases $q-f^{\prime}(\bar{u})F(\bar{u})\geq 0$ for large $\bar{u}>0$, and hence by (4.8) we have
(4.9) $\partial_{t}\bar{u}\geq\Delta\bar{u}+f(\bar{u})$
for large $\bar{u}>0$, Thus, $\bar{u}$ is a supersolution of problem (1.1) in the classical sense if $\bar{u}$ is large.
We study an intersection point of $u^{*}(x)$, which is a positive singular stationary solution of (1.1) given by Proposition 1.2, and $\bar{u}(x,t)$.
Lemma 4.2.
There exists a small $t_{0}>0$ such that $\bar{u}(r,t)$ and $u^{*}(r)$ have the first intersection point $r=r(t)\in C^{1}(0,t_{0})$ such that $r(t)\to 0$ as $t\to 0^{+}$. Specifically, there exists a positive continuous function $\delta(t)\in C(0,t_{0})$ such that
(4.10) $\begin{cases}\bar{u}(r,t)<u^{*}(r)&\text{for}\ 0<r<r(t),\\ \bar{u}(r(t),t)=u^{*}(r(t))\ \text{and}\ \bar{u}_{r}(r(t),t)>u^{*\prime}(r(t)),\\ \bar{u}(r,t)>u^{*}(r)&\text{for}\ r(t)<r<r(t)+\delta(t).\end{cases}$
Proof.
First, we prove the assertion, assuming that $q=q_{f}$ ($\varepsilon=0$). In order to find an intersection point we consider the equation $\bar{u}=u^{*}$, namely
(4.11) $F^{-1}\left[tF_{q}\left[\varphi\left(\frac{r}{\sqrt{t}},\alpha_{0}\right)\right]\right]=F^{-1}\left[\frac{r^{2}}{2N-4q_{f}}(1+\theta(r))\right].$
Here, by Proposition 3.2 we see that as $r\to 0^{+}$,
(4.12) $\theta(r)\to 0\quad\text{and}\quad r\theta^{\prime}(r)\to 0.$
Let $\eta:=r/\sqrt{t}$. Then, by (4.11) we have
$F_{q}[\varphi(\eta,\alpha_{0})]=\frac{\eta^{2}}{2N-4q_{f}}(1+\theta(r)).$
Therefore,
(4.13) $\varphi(\eta,\alpha_{0})=F_{q}^{-1}\left[\frac{\eta^{2}+\eta^{2}\theta(r)}{2N-4q_{f}}\right].$
Putting $r=\sqrt{t}\eta$ into (4.12), we have
(4.14) $\theta(\sqrt{t}\eta)\to 0\quad\text{in}\ C_{\eta,loc}(0,\infty)\quad\text{as}\ t\to 0^{+}.$
Because of Proposition 4.1, by (4.14) we see that (4.13) has the first intersection $\eta_{*}(t)$ near $\eta_{0}$ for small $t>0$ such that an intersection point does not exist on $0<\eta<\eta_{*}(t)$. We define $r(t):=\sqrt{t}\eta_{*}(t)$. By the uniqueness of the solution of equation (4.4) we see that
(4.15) $\varphi^{\prime}(\eta_{0},\alpha_{0})\neq\frac{d}{d\eta}\left.F_{q}^{-1}\left[\frac{\eta^{2}}{2N-4q_{f}}\right]\right|_{\eta=\eta_{0}}.$
Using (4.12), we have
$\frac{d}{d\eta}\left(\eta^{2}\theta(\sqrt{t}\eta)\right)$ $=2\eta\theta(\sqrt{t}\eta)+\eta(\sqrt{t}\eta)\theta^{\prime}(\sqrt{t}\eta)$
(4.16) $\to 0\quad\text{in}\ C_{\eta,loc}(0,\infty)\quad\text{as}\ t\to 0^{+}.$
It follows from (4.14) and (4.16) that the LHS of (4.13) converges to $\varphi^{*}(\eta)$ in $C^{1}_{loc}(0,\infty)$ as $t\to 0^{+}$. By (4.15) we have
$\varphi^{\prime}(\eta_{*}(t),\alpha_{0})\neq\frac{d}{d\eta}\left.F_{q}^{-1}\left[\frac{\eta^{2}}{2N-4q_{f}}(1+\theta(\sqrt{t}\eta))\right]\right|_{\eta=\eta_{*}(t)}$
for small $t>0$. Applying the implicit function theorem to the function in $(\eta,t)$ defined by
$\varphi(\eta,\alpha_{0})-F_{q}^{-1}\left[\frac{\eta^{2}}{2N-4q_{f}}(1+\theta(\sqrt{t}\eta))\right]=0,$
we see that $\eta_{*}(t)$, which is a function in $t$, is of class $C^{1}$. Thus, $r(t)=\sqrt{t}\eta_{*}(t)$ is of class $C^{1}$. Since $\eta_{*}(t)\to\eta_{0}$ as $t\to 0^{+}$, we see that $r(t)\to 0$ as $t\to 0^{+}$.
Then, it is obvious that there exists $t_{0}>0$ such that (4.11) has the first intersection point $r(t)$ satisfying (4.10) for $0<t<t_{0}$. The proof for the case $q_{f}=1$ is complete, since $q=q_{f}$.
Second, we consider the case $q_{f}>1$. Let us recall that $q=q_{f}+\varepsilon$. Let $X:=\{q\in\mathbb{R};\ q\geq 1\}$. For each compact interval $I\subset(0,\infty)$, the mappings $q\mapsto\varphi(\eta,\alpha_{0})$ and $q\mapsto F^{-1}_{q}[\eta^{2}/(2N-4q_{f})]$ are in $C(X,C^{1}_{\eta}(I))$. Since (4.15) with $q=q_{f}$ holds, we can choose a small $\varepsilon>0$ such that $q_{f}+\varepsilon<q_{S}$ and
$\varphi(\eta,\alpha_{0})=F_{q}^{-1}\left[\frac{\eta^{2}}{2N-4q_{f}}\right]$
with $q=q_{f}+\varepsilon$ has the first intersection for small $\varepsilon>0$, which can be transversal, denoted by $\tilde{\eta}_{0}$. Because of (4.14) and (4.16),
$\varphi(\eta,\alpha_{0})=F_{q}^{-1}\left[\frac{\eta^{2}+\eta^{2}\theta(\sqrt{t}\eta)}{2N-4q_{f}}\right]$
with $q=q_{f}+\varepsilon$ also has the first intersection point for small $t>0$, which can also be transversal, denoted by $\tilde{\eta}_{*}(t)$. Note that $\theta$ depends on $q_{f}$ but it does not depend on $q$. We see that $\tilde{\eta}_{*}(t)\to\tilde{\eta}_{0}$ as $t\to 0$. Let $r(t):=\sqrt{t}\tilde{\eta}_{*}(t)$. Since the first intersection point $\tilde{\eta}_{*}(t)$ is transversal, in a similar way to before we see that there exists $t_{0}>0$ such that (4.11) has the first intersection point $r(t)$ satisfying (4.10) for $0<t<t_{0}$. The proof for the case $q_{f}>1$ is complete. ∎
Now, we define a function $v$ by
(4.17) $v(x,0)=u^{*}(|x|),\quad v(x,t):=\begin{cases}\bar{u}(|x|,t)&\text{if}\ 0\leq|x|<r(t)\ \text{and}\ 0<t<t_{0},\\ u^{*}(|x|)&\text{if}\ |x|\geq r(t)\ \text{and}\ 0<t<t_{0},\end{cases}$
where $r(t)$ and $t_{0}$ are given in Lemma 4.2. By the same argument as in the proof of Proposition 1.2, we see that for every $0\leq t<t_{0}$, $v(t)\in\mathcal{L}_{ul}^{\gamma}(\mathbb{R}^{N})$ for $1\leq\gamma<\gamma^{*}$, where $\gamma_{*}$ is as in Proposition 1.2. Since $r(t)\to 0$ as $t\to 0^{+}$, we see that $u^{*}(r(t))\to\infty$ as $t\to 0^{+}$. Then, $\bar{u}(x,t)$ is large in $\{x\in\mathbb{R}^{N};\ |x|\leq r(t)\}$ for small $t>0$, and hence $\bar{u}$ becomes a supersolution in the classical sense near the origin for small $t>0$ as mentioned after (4.9).
We use the following proposition in this section.
Proposition 4.3 ([18, Corollary 3.1]).
Let $1\leq p\leq q\leq\infty$. Then there exists $C_{1}=C_{1}(N,p,q)>0$ such that
$\left\|S(t)u\right\|_{L^{q}_{ul}(\mathbb{R}^{N})}\leq C_{1}\left(t^{-\frac{N}{2}\left(\frac{1}{p}-\frac{1}{q}\right)}+1\right)\left\|u\right\|_{L^{p}_{ul}(\mathbb{R}^{N})}$
for $t>0$ and $u\in L^{p}_{ul}(\mathbb{R}^{N})$.
Then, the following lemma plays a crucial role in this paper
Lemma 4.4.
Let $v$ be defined by (4.17). Then, $v$ is a supersolution of problem (1.1) in $\mathbb{R}^{N}\times[0,t_{0})$ in the sense of Definition 1.1.
Proof.
Let $r(t)$ and $t_{0}$ be given in Lemma 4.2. Let $0<s<t_{0}$. $B_{s}$ denotes $B(0,r(s))$ for simplicity, and $B_{s}^{c}:=\mathbb{R}^{N}\setminus B_{s}$. By Green’s formula we have
(4.18) $\int_{B_{s}}G_{t-s}(x-y)\Delta\bar{u}(y)\mathop{}\!\mathrm{d}y\\ =\int_{\partial B_{s}}\left(G_{t-s}(x-y)\frac{\partial\bar{u}}{\partial{\nu_{y}}}-\bar{u}\frac{\partial}{\partial\nu_{y}}G_{t-s}(x-y)\right)\mathop{}\!\mathrm{d}S_{y}+\int_{B_{s}}\Delta_{y}G_{t-s}(x-y)\bar{u}(y)\mathop{}\!\mathrm{d}y,$
(4.19) $\int_{B_{s}^{c}}G_{t-s}(x-y)\Delta u^{*}(y)\mathop{}\!\mathrm{d}y\\ =\int_{\partial B_{s}^{c}}\left(G_{t-s}(x-y)\frac{\partial u^{*}}{\partial{\nu_{y}}}-u^{*}\frac{\partial}{\partial\nu_{y}}G_{t-s}(x-y)\right)\mathop{}\!\mathrm{d}S_{y}+\int_{B_{s}^{c}}\Delta_{y}G_{t-s}(x-y)u^{*}(y)\mathop{}\!\mathrm{d}y,$
where $\partial/\partial{\nu_{y}}$ denotes the outer unit normal derivative. Adding (4.18) and (4.19), we have
${\int_{B_{s}}G_{t-s}(x-y)\Delta v(y)\mathop{}\!\mathrm{d}y}$ ${+\int_{B^{c}_{s}}G_{t-s}(x-y)\Delta v(y)\mathop{}\!\mathrm{d}y}$
$=\int_{\partial B_{s}}G_{t-s}(x-y)\frac{\partial\bar{u}}{\partial{\nu_{y}}}\mathop{}\!\mathrm{d}S_{y}+\int_{\partial B_{s}^{c}}G_{t-s}(x-y)\frac{\partial u^{*}}{\partial{\nu_{y}}}\mathop{}\!\mathrm{d}S_{y}$
$\quad-\int_{\partial B_{s}}\bar{u}\frac{\partial}{\partial{\nu_{y}}}G_{t-s}(x-y)\mathop{}\!\mathrm{d}S_{y}-\int_{\partial B_{s}^{c}}u^{*}\frac{\partial}{\partial{\nu_{y}}}G_{t-s}(x-y)\mathop{}\!\mathrm{d}S_{y}$
$\qquad+\int_{B_{s}}\Delta_{y}G_{t-s}(x-y)\bar{u}(y)\mathop{}\!\mathrm{d}y+\int_{B_{s}^{c}}\Delta_{y}G_{t-s}(x-y)u^{*}(y)\mathop{}\!\mathrm{d}y$
(4.20) $\geq\int_{\mathbb{R}^{N}}\Delta_{y}G_{t-s}(x-y)v(y)\mathop{}\!\mathrm{d}y,$
where we used
[…]
Let $\delta>0$ be small. Integrating (4.20) over $(\delta,t-\delta)$, we have
(4.21) ${\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)\Delta v\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{B_{s}^{c}}G_{t-s}(x-y)\Delta v\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s}\\ \geq-\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}\partial_{s}G_{t-s}(x-y)v\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s,$
where we used $\Delta_{y}G_{t-s}(x-y)=-\partial_{s}G_{t-s}(x-y)$.
Let
$\mathbb{B}_{\delta}:=\left\{(y,s)\in\bigcup_{\delta<\tau<t-\delta}B_{\tau}\times\{\tau\}\right\}\quad\text{and}\quad\mathbb{B}_{\delta}^{c}:=\left\{(y,s)\in\bigcup_{\delta<\tau<t-\delta}B_{\tau}^{c}\times\{\tau\}\right\}$
Then, $\bar{u}$ and $u^{*}$ satisfy the following in the classical sense:
$\partial_{s}\bar{u}-\Delta_{y}\bar{u}=f(\bar{u})\quad\text{in}\ \mathbb{B}_{\delta}\quad\text{and}\quad\partial_{s}u^{*}-\Delta_{y}u^{*}=f(u^{*})\quad\text{in}\ \mathbb{B}_{\delta}^{c}.$
Integrating them, we have
(4.22) $\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)\bar{u}_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\\ =\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)\Delta\bar{u}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)f(\bar{u})\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s,$
(4.23) $\int_{\delta}^{t-\delta}\int_{B_{s}^{c}}G_{t-s}(x-y)u^{*}_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\\ =\int_{\delta}^{t-\delta}\int_{B_{s}^{c}}G_{t-s}(x-y)\Delta u^{*}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{B_{s}^{c}}G_{t-s}(x-y)f(u^{*})\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s.$
We consider the LHSs of (4.22) and (4.23). By Fubini’s theorem and integration by parts we have
$\int_{\delta}^{t-\delta}\int_{B_{s}}$ $G_{t-s}(x-y)\bar{u}_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s=\int_{\mathbb{B}_{\delta}}G_{t-s}(x-y)\bar{u}_{s}\mathop{}\!\mathrm{d}(y,s)$
$=\int_{\partial\mathbb{B}_{\delta}}G_{t-s}(x-y)\bar{u}\nu_{1}^{(s)}\mathop{}\!\mathrm{d}S_{{(y,s)}}-\int_{\mathbb{B}_{\delta}}\partial_{s}G_{t-s}(x-y)\bar{u}\mathop{}\!\mathrm{d}(y,s)$
(4.24) $=\int_{\partial\mathbb{B}_{\delta}}G_{t-s}(x-y)\bar{u}\nu_{1}^{(s)}\mathop{}\!\mathrm{d}S_{{(y,s)}}-\int_{\delta}^{t-\delta}\int_{B_{s}}\partial_{s}G_{t-s}(x-y)\bar{u}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s,$
$\int_{\delta}^{t-\delta}\int_{B_{s}^{c}}$ $G_{t-s}(x-y)u^{*}_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s=\int_{\mathbb{B}_{\delta}^{c}}G_{t-s}(x-y)u_{s}^{*}\mathop{}\!\mathrm{d}(y,s)$
$=\int_{\partial\mathbb{B}_{\delta}^{c}}G_{t-s}(x-y)u^{*}\nu_{2}^{(s)}\mathop{}\!\mathrm{d}S_{{(y,s)}}-\int_{\mathbb{B}_{\delta}^{c}}\partial_{s}G_{t-s}(x-y)u^{*}\mathop{}\!\mathrm{d}(y,s)$
(4.25) $=\int_{\partial\mathbb{B}_{\delta}^{c}}G_{t-s}(x-y)u^{*}\nu_{2}^{(s)}\mathop{}\!\mathrm{d}S_{{(y,s)}}-\int_{\delta}^{t-\delta}\int_{B_{s}^{c}}\partial_{s}G_{t-s}(x-y)u^{*}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s,$
where $\nu_{1}^{(s)}$ (resp. $\nu_{2}^{(s)}$) denotes the $s$ element of the outer unit normal vector on $\partial\mathbb{B}_{\delta}$ (resp. $\partial\mathbb{B}^{c}_{\delta}$) and in the integration by parts in (4.25) we used the fact that $G_{t-s}(x-y)$ is exponentially small for large $|y|$. Note that $\bigcup_{\delta<\tau<t-\delta}\partial B_{\tau}\times\{\tau\}(\subset\partial\mathbb{B}_{\delta}\cap\partial\mathbb{B}_{\delta}^{c})$ is a $C^{1}$ surface (Lemma 4.2) and hence the boundary integrals in (4.24) and (4.25) are well defined. Adding (4.24) and (4.25), we have
${\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)v_{s}}$ ${\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{B_{s}^{c}}G_{t-s}(x-y)v_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s}$
$=\left.\int_{\mathbb{R}^{N}}G_{t-s}(x-y)v\mathop{}\!\mathrm{d}y\right|_{s=t-\delta}-\left.\int_{\mathbb{R}^{N}}G_{t-s}(x-y)v\mathop{}\!\mathrm{d}y\right|_{s=\delta}$
$\qquad-\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}\partial_{s}G_{t-s}(x-y)v\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s$
(4.26) $=S(\delta)v(t-\delta)-S(t-\delta)v(\delta)-\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}\partial_{s}G_{t-s}(x-y)v\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s.$
Here we used $\nu_{1}^{(s)}+\nu_{2}^{(s)}=0$ and
$\int_{\partial\mathbb{B}_{\delta}\cap\Gamma}G_{t-s}(x-y)\bar{u}\nu_{1}^{(s)}\mathop{}\!\mathrm{d}S_{{(y,s)}}+\int_{\partial\mathbb{B}_{s}^{c}\cap\Gamma}G_{t-s}(x-y)u^{*}\nu_{2}^{(s)}\mathop{}\!\mathrm{d}S_{{(y,s)}}=0,$
where $\Gamma:=\left\{(y,s)\in\bigcup_{\delta<\tau<t-\delta}\partial B_{\tau}\times\{\tau\}\right\}$. Adding (4.22) and (4.23), we have
(4.27) ${\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)v_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{B^{c}_{s}}G_{t-s}(x-y)v_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s}\\ ={\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)\Delta v\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{B^{c}_{s}}G_{t-s}(x-y)\Delta v\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s}\\ +\int_{\delta}^{t-\delta}{\int_{\mathbb{R}^{N}}}G_{t-s}(x-y)f(v)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s.$
Using (4.21) and (4.26), by (4.27) we have
(4.28) $S(\delta)v(t-\delta)-S(t-\delta)v(\delta)-\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}\partial_{s}G_{t-s}(x-y)v\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\\ \geq-\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}\partial_{s}G_{t-s}(x-y)v\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)f(v)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s.$
By the monotone convergence theorem we have
$\lim_{\delta\to{0^{+}}}\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)f(v)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s=\int_{0}^{t}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)f(v)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s$
for a.a. $(x,y)\in\mathbb{R}^{N}\times(0,t_{0})$. Using Proposition 4.3, for $1\leq\gamma<\gamma^{*}$, we have
$\left\|S(\delta)v(t-{\delta})-v(t)\right\|_{L^{{\gamma}}_{ul}(\mathbb{R}^{N})}$
$\qquad\leq\left\|S(\delta)v(t-\delta)-S(\delta)v(t)\right\|_{L^{{\gamma}}_{ul}(\mathbb{R}^{N})}+\left\|S(\delta)v(t)-v(t)\right\|_{L^{{\gamma}}_{ul}(\mathbb{R}^{N})}$
$\qquad\leq C_{0}\left\|v(t-\delta)-v(t)\right\|_{L^{{\gamma}}_{ul}(\mathbb{R}^{N})}+\left\|S(\delta)v(t)-v(t)\right\|_{L^{{\gamma}}_{ul}(\mathbb{R}^{N})}$
$\qquad\to 0\ \text{as}\ \delta\to 0^{+},$
where $\gamma^{*}$ is as in Proposition 1.2 and we used $\lim_{\delta\to{0^{+}}}\int_{B(0,2)}|v(t-\delta)-v(t)|^{\gamma}\,\mathop{}\!\mathrm{d}x=0$ and Proposition 3.4 with $v(t)\in\mathcal{L}^{{\gamma}}_{ul}(\mathbb{R}^{N})$. Similarly, by Proposition 4.3 we have
$\left\|S(t-\delta)v(\delta)-S(t)u^{*}\right\|_{L^{{\gamma}}_{ul}(\mathbb{R}^{N})}$
$\qquad=\left\|S(t-\delta)v(\delta)-S(t-\delta)u^{*}\right\|_{L^{{\gamma}}_{ul}(\mathbb{R}^{N})}+\left\|S(t-\delta)u^{*}-S(t)u^{*}\right\|_{L^{{\gamma}}_{ul}(\mathbb{R}^{N})}$
$\qquad\leq 2C_{1}\left\|v(\delta)-u^{*}\right\|_{L^{{\gamma}}_{ul}(\mathbb{R}^{N})}+\left\|S(t-\delta)u^{*}-S(t)u^{*}\right\|_{L^{{\gamma}}_{ul}(\mathbb{R}^{N})}$
$\qquad\to 0\ \text{as}\ \delta\to 0^{+},$
where we used $\lim_{\delta\to 0}\int_{B(0,2)}|v(\delta)-u^{*}|^{\gamma}\,\mathop{}\!\mathrm{d}x=0$. We let $\delta\to 0^{+}$ in (4.28). Then, we obtain
$v(t)-S(t)u^{*}\geq\int_{0}^{t}S(t-s)f(v(s))\mathop{}\!\mathrm{d}s$
for a.a. $(x,t)\in\mathbb{R}^{N}\times(0,t_{0})$. Thus, $v(t)$ is a supersolution of problem (1.1) in $\mathbb{R}^{N}\times[0,t_{0})$ in the sense of Definition 1.1. ∎
5. Nonuniqueness (supercritical)
Lemma 5.1.
Let $\bar{w}(t)$ be a supersolution of problem (1.1) in $\mathbb{R}^{N}\times[0,T)$ in the sense of Definition 1.1, where $T>0$. If $0\leq w_{0}\leq\bar{w}(0)$ a.e. in $\mathbb{R}^{N}$, then problem (1.1) with initial data $w_{0}$ has a positive solution $w(t)$ such that $w(t)\leq\bar{w}(t)$ a.e. in $\mathbb{R}^{N}\times[0,T)$.
Proof.
We define
(5.1) $w_{1}=0,\quad w_{k}(t)=S(t)w_{0}+\int_{0}^{t}S(t-s)f(w_{k-1}(s))\mathop{}\!\mathrm{d}s$
for $k=2,3,\ldots$. By induction we can easily see that
$0=w_{1}\leq w_{2}\leq\cdots\leq w_{n}\leq\cdots\leq\bar{w}(t)<\infty$
for a.a. $(x,t)\in\mathbb{R}^{N}\times[0,T)$. For each $t>0$, the limits $\lim_{k\to\infty}w_{k}(t)$ exists for a.a. $(x,t)\in\mathbb{R}^{N}\times[0,T)$. Thus,
$w(t):=\lim_{k\to\infty}w_{k}(t)$
for a.a. $(x,t)\in\mathbb{R}^{N}\times[0,T)$. We apply the monotone convergence theorem to (5.1). Since $w(t)\leq\bar{w}(t)<\infty$ for a.a. $(x,t)\in\mathbb{R}^{N}\times[0,T)$, we obtain
$\infty>w(t)=S(t)w_{0}+\int_{0}^{t}S(t-s)f(w(s))\mathop{}\!\mathrm{d}s$
for a.a. $(x,t)\in\mathbb{R}^{N}\times[0,T)$, and hence $w$ is the solution of problem (1.1) in $\mathbb{R}^{N}\times[0,T)$ in the sense of Definition 1.1. ∎
Proof of Theorem 1.4.
We show that problem (1.1) with $u_{0}=u^{*}$ has two positive solutions. We have already shown in Lemma 1.3 that $u^{*}$ is a solution of problem (1.1) with $u_{0}=u^{*}$ in the sense of Definition 1.1. Hereafter, we show that the problem has another solution $u(t)$ such that $u(t)\in L^{\infty}_{loc}((0,t_{0}),L^{\infty}(\mathbb{R}^{N}))$.
Let $u_{0n}(x):=\min\{u^{*}(x),n\}$. Then, $u_{0n}(x)\leq v(x,0)$ a.e. in $\mathbb{R}^{N}$. Let $\zeta_{n}(t)$ be a solution of $\zeta_{n}^{\prime}(t)=f(\zeta_{n}(t))$ for $t\in(0,t_{n})$ and $\zeta_{n}(0)=n$, where $t_{n}$ is the maximal existence time of $\zeta_{n}$. Then $u_{0n}(x)\leq\zeta_{n}(0)$ in $\mathbb{R}^{N}$. Taking sufficiently large $n$ if necessary, we can assume that $t_{n}<t_{0}$, where $t_{0}$ is as in Lemma 4.4. Since $v(t)$ and $\zeta_{n}(t)$ are supersolutions in the sense of Definition 1.1 (see Lemma 4.4), we have
$\begin{split}\infty>v(t)&\geq S(t)v(0)+\int_{0}^{t}S(t-s)f(v(s))\,\mathop{}\!\mathrm{d}s\\ &\geq S(t)u_{0n}+\int_{0}^{t}S(t-s)f(\min\{v(s),\zeta_{n}(s)\})\,\mathop{}\!\mathrm{d}s\end{split}$
for a.a. $(x,t)\in\mathbb{R}^{N}\times[0,t_{0})$ and
$\begin{split}\infty>\zeta_{n}(t)&=n+\int_{0}^{t}f(\zeta_{n}(s))\,\mathop{}\!\mathrm{d}s\\ &=S(t)\zeta_{n}(0)+\int_{0}^{t}S(t-s)f(\zeta_{n}(s))\,\mathop{}\!\mathrm{d}s\\ &\geq S(t)u_{0n}+\int_{0}^{t}S(t-s)f(\min\{v(s),\zeta_{n}(s)\})\,\mathop{}\!\mathrm{d}s\end{split}$
for a.a. $(x,t)\in\mathbb{R}^{N}\times[0,t_{n})$. Noting that $\min\{v(x,t),\zeta_{n}(t)\}=v(x,t)$ for a.a. $\mathbb{R}^{N}\times[t_{n},t_{0})$, we see that these imply that $\min\{v(x,t),\zeta_{n}(t)\}$ is a supersolution of problem (1.1) in $\mathbb{R}^{N}\times[0,t_{0})$ in the sense of Definition 1.1. Thus, by Lemma 5.1 we see that problem (1.1) with initial data $u_{0n}$ has a solution $u_{n}(t)$ in $\mathbb{R}^{N}\times[0,t_{0})$ such that
$u_{n}(t)\leq v(t)\quad\text{a.e.~in}\ \mathbb{R}^{N}\times[0,t_{0})$
and
$u_{n}(t)\leq\zeta_{n}(t)\quad\text{a.e.~in}\ \mathbb{R}^{N}\times[0,t_{n}).$
Since $v\in L^{\infty}(\mathbb{R}^{N}\times(\delta,t_{0}))$ for any $\delta\in(0,t_{0})$ and $\zeta_{n}(t)<\infty$ for all $t\in[0,t_{n}/2)$, we see that $u_{n}\in L^{\infty}(\mathbb{R}^{N}\times[0,t_{0}))$, hence $u_{n}(t)$ is a classical solution of problem (1.1) in $\mathbb{R}^{N}\times[0,t_{0})$. We shall show that $u_{n}\in C(\mathbb{R}^{N}\times[0,t_{0}))$. The solution $u_{n}$ satisfies
$u_{n}(t)=S(t)u_{0n}+\int_{0}^{t}S(t-s)f(u_{n}(s))\mathop{}\!\mathrm{d}s$
for all $(x,t)\in\mathbb{R}^{N}\times[0,t_{0})$. Noting that $u_{n}(t)\in C^{2,1}(\mathbb{R}^{N}\times(0,t_{0}))\cap L^{\infty}(\mathbb{R}^{N}\times(0,t_{0}))$, we have
$\|u_{n}(t)-u_{0n}\|_{C(\mathbb{R}^{N})}\leq\|S(t)u_{0n}-u_{0n}\|_{C(\mathbb{R}^{N})}+\int_{0}^{t}\|S(t-s)f(u_{n})\|_{C(\mathbb{R}^{N})}\,\mathop{}\!\mathrm{d}s$
for all $t\in(0,t_{0})$. Since $u_{0n}\in L^{\infty}(\mathbb{R}^{N})\cap BUC(\mathbb{R}^{N})$, the first term on the RHS tends to 0 as $t\to 0^{+}$ (see e.g. [4, Section 2.3]). It suffices to consider the second term on the RHS, but by the smoothing effect of the heat kernel, it is obvious that
$\int_{0}^{t}\|S(t-s)f(u_{n})\|_{C(\mathbb{R}^{N})}\,\mathop{}\!\mathrm{d}s\leq Ctf(\|u_{n}\|_{L^{\infty}(\mathbb{R}^{N}\times(0,t_{0}))})\to 0\quad\mbox{as}\quad t\to 0^{+}.$
Thus, we see that $u_{n}\in C(\mathbb{R}^{N}\times[0,t_{0}))$. Next, we shall show that $|\nabla u_{n}|$ is bounded on $\mathbb{R}^{N}\times(0,t_{0})$. Since $u$ is smooth, we have
$\nabla u_{n}(t)=\nabla S(t)u_{0n}+\int_{0}^{t}\nabla S(t-s)f(u_{n}(s))\,\mathop{}\!\mathrm{d}s$
for all $(x,t)\in\mathbb{R}^{N}\times(0,t_{0})$. Now by integration by parts, we have
$\begin{split}&\nabla S(t)u_{0n}(x)\\ &=\nabla_{x}\left(\int_{\{u^{*}(y)\geq n\}}G_{t}(x-y)n\,\mathop{}\!\mathrm{d}y+\int_{\{u^{*}(y)<n\}}G_{t}(x-y)u^{*}(y)\,\mathop{}\!\mathrm{d}y\right)\\ &=-\int_{\partial\{u^{*}(y)\geq n\}}G_{t}(x-y)n\nu_{1}(y)\,\mathop{}\!\mathrm{d}{S_{y}}\\ &\qquad+\int_{\{u^{*}(y)<n\}}G_{t}(x-y)\nabla u^{*}(y)\,\mathop{}\!\mathrm{d}y-\int_{\partial\{u^{*}(y)<n\}}G_{t}(x-y)u^{*}(y)\nu_{2}(y)\,\mathop{}\!\mathrm{d}{S_{y}}\\ &=\int_{\{u^{*}(y)<n\}}G_{t}(x-y)\nabla u^{*}(y)\,\mathop{}\!\mathrm{d}y\end{split}$
for all $(x,t)\in\mathbb{R}^{N}\times(0,t_{0})$, where $\nu_{1}$ (resp. $\nu_{2}$) is the outer normal unit vector on $\partial\{u^{*}(y)\geq n\}$ (resp. $\partial\{u^{*}(y)<n\}$). Note that
$\partial\{u^{*}(y)\geq n\}=\partial\{u^{*}(y)<n\}=\{u^{*}(y)=n\}$
and $\nu_{1}=-\nu_{2}$ since $u^{*}$ is radial and decreasing. Therefore, we have
$\|\nabla S(t)u_{0n}\|_{L^{\infty}(\mathbb{R}^{N})}\leq C\|\nabla u^{*}\|_{L^{\infty}(\{u^{*}(x)<n\})}<\infty$
for all $t\in(0,t_{0})$. This together with the smoothing effect of the heat kernel implies that
$\begin{split}\|\nabla u_{n}(t)\|_{L^{\infty}(\mathbb{R}^{N})}&\leq C\|\nabla u^{*}\|_{L^{\infty}(\{u^{*}(x)<n\})}+Cf(\|u_{n}\|_{L^{\infty}(\mathbb{R}^{N}\times(0,t_{0}))})\int_{0}^{t}(t-s)^{-\frac{1}{2}}\,\mathop{}\!\mathrm{d}s\\ &\leq C\|\nabla u^{*}\|_{L^{\infty}(\{u^{*}(x)<n\})}+Ct^{\frac{1}{2}}f(\|u_{n}\|_{L^{\infty}(\mathbb{R}^{N}\times(0,t_{0}))})\end{split}$
for all $t\in(0,t_{0})$, which implies that $\nabla u_{n}\in L^{\infty}(\mathbb{R}^{N}\times(0,t_{0}))$.
In summary, $u_{n}(t)\in C^{2,1}(\mathbb{R}^{N}\times(0,t_{0}))\cap C(\mathbb{R}^{N}\times[0,t_{0}))$, and $|u_{n}|$ and $|\nabla u_{n}|$ are bounded on $\mathbb{R}^{N}\times(0,t_{0})$. Then, all assumptions of the comparison principle for classical solutions on $\mathbb{R}^{N}$ ([27, Proposition 52.6 in p.617]) are satisfied. We see that for $n\geq 2$,
$u_{n-1}(x,t)\leq u_{n}(x,t)\quad\text{a.e.~in}~\mathbb{R}^{N}\times[0,t_{0}).$
Therefore, there exists a limit
$u(x,t):=\lim_{n\to\infty}u_{n}(x,t)\quad\text{a.e.~in}\ \mathbb{R}^{N}\times[0,t_{0}).$
The solution $u_{n}$ satisfies
(5.2) $u_{n}(t)=S(t)u_{0n}(t)+\int_{0}^{t}S(t-s)f(u_{n}(x))\mathop{}\!\mathrm{d}s$
a.e. in $\mathbb{R}^{N}\times[0,t_{0})$. Letting $n\to\infty$ in (5.2), by the monotone convergence theorem we have
(5.3) $u(t)=S(t)u^{*}+\int_{0}^{t}S(t-s)f(u(s))\mathop{}\!\mathrm{d}s$
a.e. in $\mathbb{R}^{N}\times[0,t_{0})$. Thus, $u(t)$ is a solution in the sense of Definition 1.1. Since $u_{n}(t)\leq v(t)$, we have
$u(x,t)\leq v(x,t)\leq\bar{u}(0,t)=F^{-1}[c_{0}t]$
a.e. in $\mathbb{R}^{N}\times[0,t_{0})$, where $c_{0}:=F_{q}[\varphi(0,\alpha_{0})]$. Then, $u(x,t)\leq\min\{u^{*}(x),F^{-1}[c_{0}t]\}$ a.e. in $\mathbb{R}^{N}\times[0,t_{0})$, and hence $u(t)\in L^{\infty}_{loc}((0,t_{0}),L^{\infty}(\mathbb{R}^{N}))$. By the parabolic regularity theorem, we see that $u$ is a classical solution of problem (1.1) with $u_{0}=u^{*}$ in $\mathbb{R}^{N}\times(0,t_{0})$.
Finally, we shall prove (1.6). Since $u$ is a classical solution, $u$ satisfies (5.3) for all $(x,t)\in\mathbb{R}^{N}\times(0,t_{0})$. For $1\leq\gamma<\gamma^{*}$, we see from $u(x,t)\leq u^{*}(x)$ for a.e. $(x,t)\in\mathbb{R}^{N}\times(0,t_{0})$ that
$\|u(t)-u^{*}\|_{L^{\gamma}_{ul}(\mathbb{R}^{N})}\leq\|S(t)u^{*}-u^{*}\|_{L^{\gamma}_{ul}(\mathbb{R}^{N})}+\int_{0}^{t}\|S(t-s)f(u^{*})\|_{L^{\gamma}_{ul}(\mathbb{R}^{N})}\,\mathop{}\!\mathrm{d}s$
for all $t\in(0,t_{0})$. It follows from Propositions 1.2 and 3.4 that the first term on the RHS tends to 0 as $t\to 0^{+}$. It suffices to consider the second term on the RHS. Note that $q_{f}<q_{0}$ and $1\leq\gamma<\gamma^{*}$ imply that
$\max\left\{1,\frac{N\gamma}{N+2\gamma}\right\}<\frac{N}{2q_{f}}.$
First, let us consider the case of $1<\gamma<\gamma^{*}$. We can take a sufficiently small $\delta>0$ and $1<\beta<\gamma$ such that
$\max\left\{1,\frac{N\gamma}{N+2\gamma}\right\}<\beta<\frac{N}{2q_{f}+2\delta}<\frac{N}{2q_{f}}.$
By (3.7), we have
$\beta<\frac{N}{2q_{f}+2\delta}\implies f(u^{*})\in L^{\beta}_{ul}(\mathbb{R}^{N}).$
Furthermore,
$\frac{N\gamma}{N+2\gamma}<\beta\implies 1-\frac{N}{2}\left(\frac{1}{\beta}-\frac{1}{\gamma}\right)>0.$
Thus, we see that
$\begin{split}\int_{0}^{t}\|S(t-s)f(u^{*})\|_{L^{\gamma}_{ul}(\mathbb{R}^{N})}\,\mathop{}\!\mathrm{d}s&\leq C\|f(u^{*})\|_{L^{\beta}_{ul}(\mathbb{R}^{N})}\int_{0}^{t}\left((t-s)^{-\frac{N}{2}\left(\frac{1}{\beta}-\frac{1}{\gamma}\right)}+1\right)\,\mathop{}\!\mathrm{d}s\\ &\to 0\quad\mbox{as}\quad t\to 0^{+},\end{split}$
and obtain (1.6). Next, let us consider the case of $\gamma=1(=\beta)$. This case is much simpler. We take a sufficiently small $\delta>0$ such that $1<N/(2q_{f}+2\delta)$. Thus, we see that
$\begin{split}\int_{0}^{t}\|S(t-s)f(u^{*})\|_{L^{1}_{ul}(\mathbb{R}^{N})}\,\mathop{}\!\mathrm{d}s&\leq Ct\|f(u^{*})\|_{L^{1}_{ul}(\mathbb{R}^{N})}\to 0\quad\mbox{as}\quad t\to 0^{+},\end{split}$
and obtain (1.6). The proof is complete. ∎
6. Nonuniqueness (critical/subcritical)
In order to study the profile of self-similar solutions we consider the problem
(6.1) $\begin{cases}\displaystyle{\varphi^{\prime\prime}+\frac{N-1}{\eta}\varphi^{\prime}+\frac{1}{2}\eta\varphi^{\prime}+\frac{\varphi}{p-1}+\varphi^{p}=0},&\text{for}\ \eta>0,\\ \varphi(0)=\alpha>0,\ \varphi^{\prime}(0)=0.\end{cases}$
It was proved in [14] that $\varphi(\eta,\alpha)>0$ for $\eta\geq 0$ if $p>p_{F}$.
Proposition 6.1 ([24, Corollary 1.2]).
Let $p_{F}<p<p_{JL}$. Then the following holds:
(6.2) $\sup\{\ell>0;\ S_{\ell}\neq\emptyset\}>L,$
where $L$ is defined by (1.11) and
[…]
In Section 4 we have already seen that
$\varphi^{*}(\eta):=F_{q}^{-1}\left[\frac{\eta^{2}}{2N-4q_{f}}\right]=L\eta^{-\frac{2}{p-1}}$
is a singular solution of equation in (6.1). It follows from (6.2) that there exists $\alpha_{0}>0$ such that $\varphi(\eta,\alpha_{0})$ and $\varphi^{*}(\eta)$ has the first intersection point in $\eta>0$. Thus, Proposition 4.1 is valid for $(p_{F}<)p_{0}<p<p_{JL}$.
Proof of Theorem 1.5.
Because of (A8), we assume the existence of a singular stationary solution $u^{*}$. Since $q_{f}<q_{0}$ and $N>2$, the proof of Lemma 3.3 is valid, and hence (3.5) holds even in the case $q_{S}\leq q_{f}<q_{0}$.
Since $q_{f}<q_{0}$, two inequalities
$N-2q_{f}-2\delta>0\quad\text{and}\quad N+1-2q_{f}-2\delta>0$
hold for small $\delta>0$. Therefore, (3) holds, and the other arguments in the proof of Lemma 1.3 are valid. Thus, Lemma 1.3 holds. Specifically, $u^{*}$ is a positive radial singular stationary solution of equation in (1.1) in the sense of Definition 1.1 even in the case $q_{S}\leq q_{f}<q_{0}$.
It follows from Proposition 4.1, which is valid for $p_{0}<p<p_{JL}$, holds for $q_{S}\leq q<q_{0}$. Then, all the arguments in Section 4 are valid in the case $q_{S}\leq q_{f}<q_{0}$. In particular, $v$ defined by (4.17) is a supersolution of problem (1.1) in $\mathbb{R}^{N}\times[0,t_{0})$ in the sense of Definition 1.1.
Using $v$, we can construct a bounded solution $u(t)$ of problem (1.1) with $u_{0}=u^{*}$. The proof of the construction is the same as the proof of Theorem 1.4. In particular, $u(t)\leq u^{*}$ a.e. in $\mathbb{R}^{N}$ for $0<t<t_{0}$. Using this inequality, in the same way as Theorem 1.4 we can obtain (1.6) provided that $1\leq\gamma<\gamma^{*}$.
Now, we have two solutions, namely $u^{*}$ and $u(t)$. The proof is complete. ∎
Acknowledgement
YM was supported by JSPS KAKENHI Grant Number 24K00530.
Conflict of interest
The authors have no relevant financial or non-financial interests to disclose.
Data Availability
Data sharing is not applicable to this article as no new data were created or analyzed in this study.